# Matchings in matroids over abelian groups, II

### v3

@@ Abstract.
1Department of Mathematics, University of California, San Diego, 9500 Gilman Dr, La Jolla, CA 92093, USA.
@@ 1.1. Matchings in abelian Groups and their linear versions
Throughout this paper, we assume that $G$ is an additive abelian group, unless explicitly stated otherwise. Let $B$ be a finite subset of $G$ that excludes the neutral element 00. For any subset, $A\subset G$ with the same cardinality as $B$, a matching from $A$ to $B$ is a bijection $f:A\to B$ such that for any element $a\in A$, the sum $a+f(a)$ does not belong to $A$. The necessary conditions for a matching from $A$ to $B$ to exist are that $|A|=|B|$ and $0\notin B$. We say that $G$ has the matching property if these are sufficient conditions as well.
@@ 1.2. Matchings in Matroids
The rank of a matroid $M$, denoted by $r(M)$, is defined as $r_{M}(E(M))$. A flat in $M$ is a set $F\subset E(M)$ with the property that adjoining any new element to $F$ strictly increases its rank. A flat of rank $r(M)-1$ is called a hyperplane.
@@
    Let $M$ and $N$ be two matroids over $G$ with $r(M)=r(N)=n>0$. Let =$\mathcal{M}=\{a_{1},\ldots,a_{n}\}$ and $\mathcal{N}=\{b_{1},\ldots,b_{n}\}$ be ordered bases of $M$ and $N$, respectively. We say $\mathcal{M}$ is matched to $\mathcal{N}$ if $a_{i}+b_{i}\notin E(M)$, for all $1\leq i\leq n$.
@@
Compatibility of matchings in matroids with matchings in the vector space setting. Note that Definition 1.1 is also consistent with the notion of matchability in vector spaces over field extensions, as defined by Eliahou and Lecouvey [8]. Let $K\subset F$ be a field extension, and let $A,B\subset F$ be $n$-dimensional $K$-subspaces of $F$. Suppose $\mathcal{A}=\{a_{1},\dots,a_{n}\}$ and =$\mathcal{B}=\{b_{1},\dots,b_{n}\}$ are ordered bases of $A$ and $B$, respectively. Then $\mathcal{A}$ is said to be matched to $\mathcal{B}$ if
@@ 2.2. Matching theory
In the next two sections, we will state and prove our main results. Throughout, we may assume that $M=(E,\mathcal{I})$ and $N=(E^{\prime},\mathcal{I^{\prime}})$ are two matroids over an abelian group $(G,+)$ with neutral element 00. Any unexplained matroid terminology used here will follow Oxley [16]. We assume that $[n]=\{1,2,\cdots,n\}$ for every positive integer $n$.
@@ 3. Sparse paving matroids
[16] If $M$ is a paving matroid of rank $n\geq 2$, then its hyperplanes form a non-trivial $(n-1)$-partition of $E(M)$. That is, the intersection of any two hyperplanes in $M$ is of size at most $n-2$.
@@
    $N$ is a sparse paving matroid,
@@
The proof is straightforward in case $n=1$. So we assume that $n>1$. Since $|E(M)|=|E(N)|$ and $0\notin E(N)$, it follows from Theorem 2.3 that there exists a matching $f:E(M)\to E(N)$ in the group sense. Let $E(M)=\{a_{1},\ldots,a_{n+1}\}$. Let =$\mathcal{M}=\{a_{1},\ldots,a_{n}\}$ be a basis for $M$. Set $b_{i}=f(a_{i})$, $1\leq i\leq n$, and $\mathcal{N}=\{b_{1},\ldots,b_{n}\}$. If $\mathcal{N}$ is a basis for $N$, then $\mathcal{M}$ is matched to $\mathcal{N}$ in the matroid sense, and we are done. So suppose $\mathcal{N}$ is not a basis for $N$. Since $|E(N)|=n+1$ and $|\mathcal{N}|=n$, then $\mathcal{N}$ is a hyperplane. Define $\mathcal{N}_{i}=\left(\mathcal{N}\setminus\{b_{i}\}\right)\cup\{b_{n+1}\}$, where $\{b_{n+1}\}=E(N)\setminus\mathcal{N}$. Note that $|\mathcal{N}_{i}\cap\mathcal{N}|=n-1$. Since $N$ is a sparse paving matroid, by Theorem 3.1 $\mathcal{N}_{i}$’s are not hyperplanes. This together with the fact that $r_{N}(\mathcal{N}_{i})\geq n-1$ implies that each $\mathcal{N}_{i}$ is a basis. We now claim that there exists $i\in[n]$ such that $a_{i}+b_{n+1}\notin E(M)$. If not, we have
@@
    $N$ is a sparse paving matroid,
@@
In general, within matching theory—even in the group setting, where the structure is relatively simpler—the existence of a total order compatible with the group operation is essential. For instance, in [6], which addresses the existence of a specific family of matchings called acyclic matchings in groups, the proof relies heavily on the presence of a well-behaved total order on $\mathbb{Z}^{n}$. Unfortunately, such a feature is available only in torsion-free abelian groups [14]. Since our goal is to establish results in a more general setting, we instead employ the rectification principle Theorem 2.2, which ensures the existence of a suitably ”nice” total order for sufficiently small subsets of the group. This is formalized through (2), which we use to guarantee such an order.
@@ 4. Panhandle matroids and Schubert matroids over Abelian Groups
$[m]_{(a,G)}=\{a,2a,\dots,ma\},$
where $ma=a+\underbrace{\cdots}_{m\text{ times}}+a$.
Let $n\leq s<m$ be nonnegative integers. The extended panhandle matroid $\mathcal{P}_{n,s,m}(a,G)$ is the rank $n$ matroid on the ground set $[m]_{(a,G)}$ with basis system
@@
We abbreviate $[m]_{(a,G)}$, $\mathcal{B}_{(n,s,m)}(a,G)$, and $\mathcal{P}_{n,s,m}(a,G)$ to $[m]_{a}$, $\mathcal{B}_{(n,s,m)}(a)$, and $\mathcal{P}_{n,s,m}(a)$ when $G$ is clear from the context.
@@
$[5]_{\left((2,-1,0),\mathbb{Z}^{3}\right)}=\left\{(2,-1,0),\,(4,-2,0),\,(6,-3,0),\,(8,-4,0),\,(10,-5,0)\right\}.$
@@
[…]
@@
In what follows, we define Schubert matroids over abelian groups. Similar to the panhandle matroid case, we assume that $G$ (or small enough subsets of $G$) is equipped with a total ordering $\preceq$ that is compatible with its structure. In particular, for a group containing nonzero torsion elements, we assume that $m$ and $n$ are two positive integers with $n\leq m<\lceil\log_{2}(p(G))\rceil$. Let $a\in G^{+}$ and $S$ be an $n$-subset of $[m]_{a}$.
The extended Schubert matroid: $SM_{m}(a,G,S)$ is the rank $n$ matroid whose ground set is $[m]_{a}$ and whose basis system is
@@
$[5]_{\left((2,-1,0),\mathbb{Z}^{3}\right)}=\left\{(2,-1,0),\,(4,-2,0),\,(6,-3,0),\,(8,-4,0),\,(10,-5,0)\right\}$
@@
=$\mathcal{B}=\left\{T\subseteq[5]_{(2,-1,0)}\mid T\preceq\left\{(2,-1,0),(4,-2,0),(10,-5,0)\right\}\right\}$
[…]
@@
First, assume that $M$ is matched to $\mathcal{P}_{n,s^{\prime},m}(a)$. From the structure of $M$, it follows that the $n$-subset =$\mathcal{M}=\{a,2a,\dots,na\}$ of $[m]_{a}$ is a basis for $M$. Consequently, it must be matched to a basis $\mathcal{N}=\{b_{1},b_{2},\dots,b_{n}\}$ of $\mathcal{P}_{n,s^{\prime},m}(a)$, where $b_{1}\prec b_{2}\prec\dots\prec b_{n}$.
@@
Conversely, assume that $s^{\prime}\neq m-1$. Thus, we have $s^{\prime}<m-1$. Consider the basis =$\mathcal{M}=\{a,2a,\dots,na\}$ of $M$. We claim that $\mathcal{M}$ cannot be matched to any basis of $\mathcal{P}_{n,s^{\prime},m}(a)$. Indeed, if $\mathcal{M}$ is matched to a basis $\mathcal{N}=\{b_{1},\dots,b_{n}\}$ of $\mathcal{P}_{n,s^{\prime},m}(a)$ with $b_{1}\prec b_{2}\prec\dots\prec b_{n}$, then it follows that $a$ must be mapped to $b_{n}=ma$.
@@
Let $G$ be an abelian group, and let $n\leq s<m$ be positive integers with $m$ sufficiently small. Let $a\in G$ be a nonzero element. Consider two subsets $S$ and $S^{\prime}$ of $[m]_{a}$ such that $|S|=|S^{\prime}|=n$. Assume that $M=SM_{m}(a,S)$ or $M=\mathcal{P}_{n,s,m}(a)$. Then, $M$ is matched to $SM_{m}(a,S^{\prime})$ if and only if $SM_{m}(a,S^{\prime})\cong U_{n,m}$.
@@
First, assume that $M$ is matched to $SM_{m}(a,S^{\prime})$. Without loss of generality, we can assume that $a$ is positive. Let $S^{\prime}=\{b_{1},\dots,b_{n}\}$, where $b_{1}\prec\dots\prec b_{n}$. Consider the $n$-subset =$\mathcal{M}=\{a,2a,\dots,na\}$ of $[m]_{a}$. By the definition of $\mathcal{P}_{n,s,m}(a)$ (or $SM_{m}(a,S)$), $\mathcal{M}$ is a basis for $M$. Therefore, it must be matched to a basis $\mathcal{N}=\{a_{1},\dots,a_{n}\}$ of $SM_{m}(a,S^{\prime})$.
@@
Conversely, assume that $SM_{m}(a,S^{\prime})\cong U_{n,m}$. Since $0\notin[m]_{a}$, it follows from Theorem 2.4 that $[m]_{a}$ is matched to itself in the group sense. That is, there exists a bijection $f:[m]_{a}\rightarrow[m]_{a}$ such that $x+f(x)\notin[m]_{a}$ for any $x\in[m]_{a}$. Let =$\mathcal{M}=\{a_{1},\dots,a_{n}\}$ be a basis for $M$, set $b_{i}=f(a_{i})$ for $i\in[n]$ and $\mathcal{N}=\{b_{1},\dots,b_{n}\}$. Since $SM_{m}(a,S^{\prime})$ is a uniform matroid with rank $n$, $\mathcal{N}$ is a basis for $SM_{m}(a,S^{\prime})$. $\mathcal{M}$ is matched to $\mathcal{N}$ in the matroid sense. ∎

### v4

@@ Abstract.
1Department of Mathematics, Clayton State University, 2000 Clayton State Boulevard, Morrow, GA 30260, USA.
2Department of Mathematics, University of California, San Diego, 9500 Gilman Dr, La Jolla, CA 92093, USA.
@@ 1.1. Matchings in abelian Groups and their linear versions
Throughout this paper, we assume that $G$ is an additive abelian group, unless explicitly stated otherwise. Let $B$ be a finite subset of $G$ that excludes the neutral element 0. For any subset, $A\subset G$ with the same cardinality as $B$, a matching from $A$ to $B$ is a bijection $f:A\to B$ such that for any element $a\in A$, the sum $a+f(a)$ does not belong to $A$. The necessary conditions for a matching from $A$ to $B$ to exist are that $|A|=|B|$ and $0\notin B$. We say that $G$ has the matching property if these are sufficient conditions as well.
@@ 1.2. Matchings in Matroids
The rank of a matroid $M$, denoted by $r(M)$, is defined as $r_{M}(E(M))$. A flat in $M$ is a set $F\subset E(M)$ with the property that adjoining any new element to $F$ strictly increases its rank. A flat of rank r(M)1r(M)-1 is called a hyperplane.
@@
    Let $M$ and $N$ be two matroids over $G$ with $r(M)=r(N)=n>0$. Let $\mathcal{M}=\{a_{1},\ldots,a_{n}\}$ and $\mathcal{N}=\{b_{1},\ldots,b_{n}\}$ be ordered bases of $M$ and $N$, respectively. We say $\mathcal{M}$ is matched to $\mathcal{N}$ if $a_{i}+b_{i}\notin E(M)$, for all $1\leq i\leq n$.
@@
Compatibility of matchings in matroids with matchings in the vector space setting. Note that Definition 1.1 is also consistent with the notion of matchability in vector spaces over field extensions, as defined by Eliahou and Lecouvey [8]. Let $K\subset F$ be a field extension, and let $A,B\subset F$ be $n$-dimensional $K$-subspaces of $F$. Suppose $\mathcal{A}=\{a_{1},\dots,a_{n}\}$ and $\mathcal{B}=\{b_{1},\dots,b_{n}\}$ are ordered bases of $A$ and $B$, respectively. Then $\mathcal{A}$ is said to be matched to $\mathcal{B}$ if
@@ 2.2. Matching theory
In the next two sections, we will state and prove our main results. Throughout, we may assume that $M=(E,\mathcal{I})$ and $N=(E^{\prime},\mathcal{I^{\prime}})$ are two matroids over an abelian group $(G,+)$ with neutral element 0. Any unexplained matroid terminology used here will follow Oxley [16]. We assume that $n]=\{1,2,\cdots,n\}$ for every positive integer $n$.
@@ 3. Sparse paving matroids
[16] If $M$ is a paving matroid of rank $n\geq 2$, then its hyperplanes form a non-trivial (n1)(n-1)-partition of $E(M)$. That is, the intersection of any two hyperplanes in $M$ is of size at most n2n-2.
@@
    NN is a sparse paving matroid,
@@
The proof is straightforward in case $n=1$. So we assume that $n>1$. Since $|E(M)|=|E(N)|$ and $0\notin E(N)$, it follows from Theorem 2.3 that there exists a matching $f:E(M)\to E(N)$ in the group sense. Let $E(M)=\{a_{1},\ldots,a_{n+1}\}$. Let $\mathcal{M}=\{a_{1},\ldots,a_{n}\}$ be a basis for $M$. Set $b_{i}=f(a_{i})$, $1\leq i\leq n$, and $\mathcal{N}=\{b_{1},\ldots,b_{n}\}$. If $\mathcal{N}$ is a basis for $N$, then $\mathcal{M}$ is matched to $\mathcal{N}$ in the matroid sense, and we are done. So suppose $\mathcal{N}$ is not a basis for $N$. Since $|E(N)|=n+1$ and $|\mathcal{N}|=n$, then $\mathcal{N}$ is a hyperplane. Define $\mathcal{N}_{i}=\left(\mathcal{N}\setminus\{b_{i}\}\right)\cup\{b_{n+1}\}$, where $\{b_{n+1}\}=E(N)\setminus\mathcal{N}$. Note that $|\mathcal{N}_{i}\cap\mathcal{N}|=n-1$. Since $N$ is a sparse paving matroid, by Theorem 3.1 $\mathcal{N}_{i}$’s are not hyperplanes. This together with the fact that $r_{N}(\mathcal{N}_{i})\geq n-1$ implies that each $\mathcal{N}_{i}$ is a basis. We now claim that there exists $i\in[n]$ such that $a_{i}+b_{n+1}\notin E(M)$. If not, we have
@@
    NN is a sparse paving matroid,
@@
In general, within matching theory—even in the group setting, where the structure is relatively simpler—the existence of a total order compatible with the group operation is essential. For instance, in [6], which addresses the existence of a specific family of matchings called acyclic matchings in groups, the proof relies heavily on the presence of a well-behaved total order on $\mathbb{Z}^{n}$. Unfortunately, such a feature is available only in torsion-free abelian groups [14]. Since our goal is to establish results in a more general setting, we instead employ the rectification principle Theorem 2.2, which ensures the existence of a suitably nice total order for sufficiently small subsets of the group. This is formalized through (2), which we use to guarantee such an order.
@@ 4. Panhandle matroids and Schubert matroids over Abelian Groups
$m]_{(a,G)}=\{a,2a,\dots,ma\},$
where ma=a+m $a+\underbrace{\cdots}_{m\text{ times}}+a$.
Let $n\leq s<m$ be nonnegative integers. The extended panhandle matroid $\mathcal{P}_{n,s,m}(a,G)$ is the rank $n$ matroid on the ground set $m]_{(a,G)}$ with basis system
@@
We abbreviate $m]_{(a,G)}$, $\mathcal{B}_{(n,s,m)}(a,G)$, and $\mathcal{P}_{n,s,m}(a,G)$ to $m]_{a}$, $\mathcal{B}_{(n,s,m)}(a)$, and $\mathcal{P}_{n,s,m}(a)$ when $G$ is clear from the context.
@@
$5]_{\left((2,-1,0),\mathbb{Z}^{3}\right)}=\left\{(2,-1,0),\,(4,-2,0),\,(6,-3,0),\,(8,-4,0),\,(10,-5,0)\right\}.$
@@
$\left\{\begin{aligned} &\left\{(2,-1,0),\,(4,-2,0),\,(6,-3,0)\right\},\\ &\left\{(2,-1,0),\,(6,-3,0),\,(8,-4,0)\right\},\\ &\left\{(2,-1,0),\,(4,-2,0),\,(8,-4,0)\right\},\\ &\left\{(4,-2,0),\,(6,-3,0),\,(8,-4,0)\right\},\\ &\left\{(2,-1,0),\,(4,-2,0),\,(10,-5,0)\right\},\\ &\left\{(2,-1,0),\,(6,-3,0),\,(10,-5,0)\right\},\\ &\left\{(2,-1,0),\,(8,-4,0),\,(10,-5,0)\right\},\\ &\left\{(4,-2,0),\,(6,-3,0),\,(10,-5,0)\right\},\\ &\left\{(4,-2,0),\,(8,-4,0),\,(10,-5,0)\right\},\\ &\left\{(6,-3,0),\,(8,-4,0),\,(10,-5,0)\right\}\end{aligned}\right\}.$
@@
In what follows, we define Schubert matroids over abelian groups. Similar to the panhandle matroid case, we assume that $G$ (or small enough subsets of $G$) is equipped with a total ordering $\preceq$ that is compatible with its structure. In particular, for a group containing nonzero torsion elements, we assume that $m$ and $n$ are two positive integers with $n\leq m<\lceil\log_{2}(p(G))\rceil$. Let $a\in G^{+}$ and $S$ be an $n$-subset of $m]_{a}$.
The extended Schubert matroid: $SM_{m}(a,G,S)$ is the rank $n$ matroid whose ground set is $m]_{a}$ and whose basis system is
@@
$5]_{\left((2,-1,0),\mathbb{Z}^{3}\right)}=\left\{(2,-1,0),\,(4,-2,0),\,(6,-3,0),\,(8,-4,0),\,(10,-5,0)\right\}$
@@
$\mathcal{B}=\left\{T\subseteq[5]_{(2,-1,0)}\mid T\preceq\left\{(2,-1,0),(4,-2,0),(10,-5,0)\right\}\right\}$
$\left\{\begin{aligned} &\left\{(2,-1,0),\,(4,-2,0),\,(6,-3,0)\right\},\\ &\left\{(2,-1,0),\,(4,-2,0),\,(8,-4,0)\right\},\\ &\left\{(2,-1,0),\,(4,-2,0),\,(10,-5,0)\right\}\end{aligned}\right\}.$
@@
First, assume that $M$ is matched to $\mathcal{P}_{n,s^{\prime},m}(a)$. From the structure of $M$, it follows that the $n$-subset $\mathcal{M}=\{a,2a,\dots,na\}$ of $m]_{a}$ is a basis for $M$. Consequently, it must be matched to a basis $\mathcal{N}=\{b_{1},b_{2},\dots,b_{n}\}$ of $\mathcal{P}_{n,s^{\prime},m}(a)$, where $b_{1}\prec b_{2}\prec\dots\prec b_{n}$.
@@
Conversely, assume that $s^{\prime}\neq m-1$. Thus, we have $s^{\prime}<m-1$. Consider the basis $\mathcal{M}=\{a,2a,\dots,na\}$ of $M$. We claim that $\mathcal{M}$ cannot be matched to any basis of $\mathcal{P}_{n,s^{\prime},m}(a)$. Indeed, if $\mathcal{M}$ is matched to a basis $\mathcal{N}=\{b_{1},\dots,b_{n}\}$ of $\mathcal{P}_{n,s^{\prime},m}(a)$ with $b_{1}\prec b_{2}\prec\dots\prec b_{n}$, then it follows that $a$ must be mapped to $b_{n}=ma$.
@@
Let $G$ be an abelian group, and let $n\leq s<m$ be positive integers with $m$ sufficiently small. Let $a\in G$ be a nonzero element. Consider two subsets $S$ and $S^{\prime}$ of $m]_{a}$ such that $|S|=|S^{\prime}|=n$. Assume that $M=SM_{m}(a,S)$ or $M=\mathcal{P}_{n,s,m}(a)$. Then, $M$ is matched to $SM_{m}(a,S^{\prime})$ if and only if $SM_{m}(a,S^{\prime})\cong U_{n,m}$.
@@
First, assume that $M$ is matched to $SM_{m}(a,S^{\prime})$. Without loss of generality, we can assume that $a$ is positive. Let $S^{\prime}=\{b_{1},\dots,b_{n}\}$, where $b_{1}\prec\dots\prec b_{n}$. Consider the $n$-subset $\mathcal{M}=\{a,2a,\dots,na\}$ of $m]_{a}$. By the definition of $\mathcal{P}_{n,s,m}(a)$ (or $SM_{m}(a,S)$), $\mathcal{M}$ is a basis for $M$. Therefore, it must be matched to a basis $\mathcal{N}=\{a_{1},\dots,a_{n}\}$ of $SM_{m}(a,S^{\prime})$.
@@
Conversely, assume that $SM_{m}(a,S^{\prime})\cong U_{n,m}$. Since $0\notin[m]_{a}$, it follows from Theorem 2.4 that $m]_{a}$ is matched to itself in the group sense. That is, there exists a bijection $f:[m]_{a}\rightarrow[m]_{a}$ such that $x+f(x)\notin[m]_{a}$ for any $x\in[m]_{a}$. Let $\mathcal{M}=\{a_{1},\dots,a_{n}\}$ be a basis for $M$, set $b_{i}=f(a_{i})$ for $i\in[n]$ and $\mathcal{N}=\{b_{1},\dots,b_{n}\}$. Since $SM_{m}(a,S^{\prime})$ is a uniform matroid with rank $n$, $\mathcal{N}$ is a basis for $SM_{m}(a,S^{\prime})$. $\mathcal{M}$ is matched to $\mathcal{N}$ in the matroid sense. ∎